# Empirical Evaluation of Generalizable Automated Program Repair with Large Language Models

### v1

@@ Abstract
Automated Program Repair (APR) proposes bug fixes to aid developers in maintaining software. The state of the art in this domain focuses on using LLMs, leveraging their strong capabilities to comprehend specifications in natural language and to generate program code. Recent works have shown that LLMs can be used to generate repairs. However, despite the APR community’s research achievements and several industry deployments in the last decade, APR still lacks the capabilities to generalize broadly. In this work, we present an intensive empirical evaluation of LLMs’ capabilities in generating patches. We evaluate a diverse set of 13 recent models, including open ones (e.g., Llama 3.3, Qwen 2.5 Coder, and DeepSeek R1 (dist.)) as well as closed ones (e.g., o3-mini, GPT-4o, Claude 3.7 Sonnet, Gemini 2.0 Flash). In particular, we explore language-agnostic repair capabilities by utilizing benchmarks for Java (e.g., Defects4J), JavaScript (e.g., BugsJS), Python (e.g., BugsInPy), and PHP (e.g., BugsPHP). Besides the generalization between different languages and levels of patch complexity, we also investigate the effects of fault localization as a preprocessing step and compare the progress for open vs closed models.
Our evaluation represents a snapshot of the current repair capabilities of the latest LLMs. Key results include that: (1) Different LLMs tend to perform best for different languages, which makes it hard to develop cross-platform repair techniques with single LLMs. (2) The combinations of models add value with respect to uniquely fixed bugs, so a committee of expert models should be considered. (3) Under realistic assumptions of imperfect fault localization, we observe significant drops in accuracy from the usual practice of using perfect fault localization. Our findings and insights will help both researchers and practitioners develop reliable and generalizable APR techniques and evaluate them in realistic and fair environments.
Index Terms:
automated program repair, large language models, empirical evaluationI Introduction
@@
Automated Program Repair (APR) [1, 2] has gained significant attention in the last few years. It attempts to repair software bugs automatically, and hence, promises great support for software practitioners in maintaining code. APR research in the last decade has focused on search-based, semantics-based, template-based, and learning-based methods. Search-based techniques [3] apply evolutionary code mutations to change the program to pass a test suite. Semantics-based techniques [4] use a test suite to extract specifications as constraints via, e.g., symbolic execution. These constraints then drive a program synthesis to generate a repair, e.g., expressions or whole statements. Template-based techniques [5] use pre-defined patch templates that are either manually crafted for a specific domain or mined from large code repositories. Finally, learning-based techniques [6] attempt to directly learn how to generate patches based on large sets of historic fixes.
@@
Despite these research achievements in APR and existing industrial deployments [7, 8, 9], there are still issues with generalizing and benchmark overfitting. The first international competition on APR [10] in 2024 using the Cerberus framework [11] has shown that current techniques are not yet able to reliably provide software patches, i.e., they failed to work in realistic settings and generalize beyond the community benchmarks. For example, in the Functional Error track for Java, the best tool was able to correctly fix only out of previously unseen bugs (i.e., 10% fix rate) and many techniques were unable to provide any correct patches. The benchmarks had been curated with new and unforeseen repair tasks, and the tools had been evaluated in a uniform environment enforcing strict resource constraints.
@@
Machine Learning (ML) techniques, specifically Large Language Models (LLM), play a significant role in the SE and APR community’s most recent research efforts. Due to their pre-training on large amounts of open-source code and due to their enormous capacity, LLMs have demonstrated problem-solving skills over a wide range of tasks [12]. In APR, recent works [13, 14] have shown that LLMs can be used to generate repairs and build SE agents [15, 16, 17], which attempt to mimic the work of software engineers. Nevertheless, most of the current LLM-based APR approaches use LLMs in a rather ad-hoc strategy (e.g., [13, 18, 19, 20]), and while some existing works study the usage of LLMs in a more systematic way (e.g., [21, 22, 23]), they all fall short on at least one of the following points: (i) the investigation is based on one specific LLM only, (ii) no/limited exploration of various prompting strategies/designs is conducted, in particular regarding repair ingredients (like information about failing test cases, bug reports/descriptions, error messages, etc.), (iii) a perfect Fault Localization (FL) is assumed to be given, which represents an unrealistic assumption in a practical setting, and (iv) only one specific benchmark with one programming language is covered, e.g., Defects4J [24].
@@
To address the above issues and to provide a comprehensive overview of the current research landscape of APR, this work presents a systematic empirical evaluation with state-of-the-art closed and open LLMs on multiple benchmarks across programming languages. Hereby, we refer to open models as those with publicly available weights111Note that, according to our definition, open models may not necessarily come with open-source licenses, and that – unlike their weights – their training code and/or data may remain undisclosed., whereas closed LLMs are proprietary models whose weights, training data, and model architectures have not been publicly released. Particularly, we compare closed models like Claude 3.7 Sonnet, Gemini 1.5 Pro, OpenAI o3-mini, and OpenAI GPT-4o with open ones like Llama 3.3, CodeLlama, and DeepSeek Coder on benchmarks for Java (Defects4J [24]), JavaScript (BugsJS [25]), Python (BugsInPy [26]), and PHP (BugsPHP [27]).
@@
Based on our experiments, current LLMs do not generalize well across programming languages for the task of program repair. In particular, we observed that different LLMs perform well on different benchmarks. Similar to others [28], we observed that including information about the failing test case significantly boosted the repair performance for all models. We found LLMs to show good robustness when repairing bugs of higher complexity. However, notably, our comparison between perfect and automated FL techniques showed a drastic decline in repair capabilities if the buggy function cannot reliably be identified. Overall, our findings provide fundamental and comprehensive understanding of the current capabilities of LLMs with respect to the various aspects of APR, which is beneficial for our community (e.g., for designing future research efforts) and for practitioners (e.g., for selecting appropriate LLMs for their repair use case). In summary, we make the following core contributions:
@@
To the future APR researcher, our study shows that (1) fault localization is a potential bottleneck for APR pipelines, which suggests a strong focus on this aspect. (2) Overall, we found APR accuracy to generalize poorly, i.e., different models react differently when varying programming language or prompt information. We conclude that APR pipelines need to be benchmarked carefully and broadly. To support such a holistic evaluation of future progress in APR, we open-source our experimental setup and results for other researchers.
II Background and Related Work
II-A LLM based Program Repair
Numerous APR techniques have been studied in the last decade, ranging from the traditional search-based repair [3], constraint-based repair [4], template-based repair [5], and learning-based repair [29]. More recently, with the advancement of LLMs for code generation, various techniques [13, 14, 6] have been explored by the research community to utilize LLMs for the program repair problem. LLM-based repair techniques transform the program repair problem into a code generation problem. These techniques can be broadly classified as model training [22, 18], zero-shot learning [13] and few-shot learning [30, 8, 31]. Silva et al. [22] explored how different input and output representations that yield optimized results using LLMs for program repair tasks. Xia et al. [14] studied the effect of test diagnostic feedback for LLM-driven program repair. Lin et al. [32] studied the impact of repair granularities encountered in program repair tasks. More specifically, how expression-level repair to token-level repair can differ for different LLMs.
@@ II-A LLM based Program Repair
II-B Empirical Studies of Program Repair
@@
Several empirical studies have been conducted to evaluate program repair techniques in their efficacy [33], performance [34], trust [35], and overfitting [36, 37] aspects. Ye et al. [37] conducted an empirical study using 10 pre-LLM program repair tools for Java programs and found overfitting to the respective test suites to be an issue. Similarly, Durieux et al. [36] conducted a large-scale evaluation using 11 APR tools to fix 2,141 Java programs in five different repair benchmarks. Their experiments showed that repair tools do not generalize across benchmarks, highlighting the problem of benchmark overfitting in repair techniques. Liu et al. [33] conducted another study on the efficiency of repair tools using 16 Java repair tools. Their study highlighted the performance inefficiencies in repair tools, which predominantly generated a large number of incorrect patches despite their overall repair capabilities. Noller et al. [35] performed a small-scale experiment using five program repair tools and 60 C programs. Their findings reported that different experiment configurations lead to different results, highlighting the importance of fair comparison under a uniform setup. A similar study was conducted by Zhong et al. [34] for learning-based repair tools, highlighting the importance of a unified experimental setup with respect to the training data and evaluation metrics used for comparisons.
@@
More recently, Xia et al. [38] conducted an extensive evaluation using 9 LLMs and five repair benchmarks covering multiple languages under different repair configurations. Their findings highlighted the benefits of using LLMs for APR, with respect to the quality of the generated patches and the performance improvements compared to non-LLM-based APR tools. Several other studies [23, 21] investigated fine-tuned coding models for program repair using multiple benchmarks, showing they can outperform non-LLM-based repair tools. Ouyang et al. [39] studied 11 repair tools including traditional techniques, NMT-based, and LLM-based repair tools. Their findings further emphasized the benefits of using LLM-based repair tools, which showed less susceptibility to overfitting and better generalization across benchmarks. Unlike existing work, in this study we perform an extensive analysis using the most recent 13 LLMs including the most advanced reasoning models across various experiment configurations. We focus on the generalizability of LLMs across different languages, different repair configurations and different patch complexities. We also revisit the impact of accurate fault localization on program repair with respect to the usage of the most advanced reasoning LLMs.
III Study Methodology
@@
The primary objective of this work is to explore the feasibility of APR using current LLMs. We first discuss the four research questions addressed in our study (Section III-A), followed by the LLMs we evaluate (Section III-B) and the benchmark datasets in different programming languages (Section III-C). We outline further technical details – the evaluation protocol, technical setup and prompt design – in Sections III-D to III-E.
III-A Research Questions
@@
To answer this question, we apply LLM-based APR on repair benchmarks for Java, Python, PHP, and JavaScript and assess the quality of the generated patches. As part of this question, we also investigate whether one model appears superior across all programming languages, or whether different models work best on specific languages (which would suggest a language-specific model selection). To investigate this question, we use two prompt designs (see Section III-E); (1) a basic prompt including only the code passage to be repaired, and (2) an extended prompt which also includes additional information from failing tests’ error messages.
@@
Traditionally, APR pipelines feature a fault localization (FL) step (in which a faulty code subsegment is identified) and a repair step (in which the subsegment is fixed). LLMs could either follow this approach and take localization hints from external systems such as FLACOCO [40] using spectrum-based fault localization [41], or operate in an end-to-end fashion, fixing a given chunk of code without explicitly localizing faulty subsegments. So far, there are no conclusive research results as to what extent LLMs benefit from localization, with a lot of research assuming perfect localization to be given [13, 14, 20, 21, 22, 23], others focusing on end-to-end operation [42], and only few recent comparisons of both with inconclusive results so far [22, 43]. To this end, we investigate whether current LLMs benefit from localization signals at code line granularity.
@@
In our last question, we focus on the differences in APR accuracy between open LLMs and closed LLMs. While closed models have been found to offer superior performance on the majority of natural language understanding (NLU) tasks in the past (e.g., [12]), open LLMs offer the benefit of being operational locally, thus allowing for better reproducibility of results and supporting technological sovereignty and democratization. Therefore, we track the progress of both categories of LLMs over time, to investigate whether open models are catching up with their closed counterparts.
III-B Models
@@ III-C Benchmarks
Finally, the benchmarks should offer a sufficient number of bugs that are still reproducible today. As Zhu et al. [44] have shown, reproducibility is often a challenge in executable software defects datasets, as legacy versions of software artifacts are prone to breakage due to missing/deprecated dependencies or external resources.
Similar to prior studies [13, 38, 39], we identified bugs of comparable complexity within the benchmarks to enable fair comparisons across programming languages. To do so, we evaluated single-function fixes within the benchmarks, meaning we considered only those bugs that could be resolved by modifications within a single function. Unlike many previous studies, however, we did not restrict our analysis to single-line[5, 45] or single-hunk fixes[13], that is, bugs that can be resolved by modifying a single contiguous piece of code.
@@
Defects4J v3.0 [24]
This is a curated collection of reproducible bugs with a supporting infrastructure and has been widely adopted in software engineering research. Most prior studies on LLM-based APR were performed on Defects4J [24], thus enabling a direct comparison with prior work. It contains 854 bugs from 17 open-source Java projects. For our first research question, we use the 264 single-hunk bugs which we could reproduce.
BugsInPy [26]
This is a benchmark dataset comprising 493 real bugs from 17 open-source Python projects across diverse domains, including machine learning, developer tools, scientific computing, and web frameworks. Unlike Defects4J, which distinguishes between active and deprecated bugs (i.e. those no longer reproducible with current Java versions), BugsInPy does not track the reproducibility status of its bugs. Aguilar et al. [46] evaluated the reproducibility of the dataset three years after its initial release and found that only of the expected results could be reproduced in the unmodified BugsInPy dataset. Among these 337 reproducible bugs, 82 were single-hunk bugs, which we used in our first experiment.
BugsJS [25]
@@
BugsPHP [27]
BugsPHP is a benchmark dataset of PHP bugs on real-world applications consisting of a training and test dataset. We use BugsPHP’s test split, which contains 513 bug-fixing commits from 15 popular open-source PHP projects covering a time period of Jan 2020 – Mar 2023. Each commit is equipped with at least one relevant failing test case. For our initial experiment, we identified 129 reproducible single-hunk bugs.
| Benchmark | all bugs | single-file | s.-function | s.-hunk | single-line |

### v2

# Exploring Generalizable Automated Program Repair with Large Language Models

@@ Abstract.
Automated Program Repair (APR) proposes bug fixes to aid developers in maintaining software. The state of the art in this domain focuses on LLMs, leveraging their strong capabilities to comprehend specifications in natural language and to generate program code. However, despite the APR community’s research achievements and industry deployments, APR still cannot generalize broadly. In this work, we present an intensive empirical evaluation of LLMs’ capabilities in APR. We evaluate a diverse set of 13 recent open and closed models. In particular, we explore language-agnostic repair by utilizing benchmarks for Java, JavaScript, Python, and PHP. Besides the generalization across languages and levels of patch complexity, we also investigate the effects of fault localization (FL). Our key results include: (1) Different LLMs tend to perform best for different languages, which makes it hard to develop cross-platform, single-LLM repair techniques. (2) Combining models by pooling repairs adds value with respect to uniquely fixed bugs, so a committee of expert models should be considered. (3) Under realistic assumptions of imperfect FL, we observe significant drops in accuracy from the usual practice of using perfect FL. Our insights will help develop reliable and generalizable APR techniques and evaluate them in realistic and fair environments.
1. Introduction
@@
Automated Program Repair (APR) (Le Goues et al., 2021; Gao et al., 2022) has gained significant attention in the last few years. It attempts to repair software bugs automatically, and hence, promises great support for software practitioners in maintaining code. APR research in the last decade has focused on search-based, semantics-based, template-based, and learning-based methods. Search-based techniques (Le Goues et al., 2012) apply evolutionary code mutations to change the program to pass a test suite. Semantics-based techniques (Nguyen et al., 2013) use a test suite to extract specifications as constraints via, e.g., symbolic execution. These constraints then drive a program synthesis to generate a repair, e.g., expressions or whole statements. Template-based techniques (Liu et al., 2019b) use pre-defined patch templates that are either manually crafted for a specific domain or mined from large code repositories. Finally, learning-based techniques (Ye and Monperrus, 2024) attempt to directly learn how to generate patches based on large sets of historic fixes.
@@
Despite these research achievements in APR and existing industrial deployments (Williams et al., 2024; Jin et al., 2023; Marginean et al., 2019), there are still issues with generalizing and benchmark overfitting. The first international competition on APR (Shariffdeen et al., 2024) in 2024 using the Cerberus framework (Shariffdeen et al., 2023) has shown that current techniques are not yet able to reliably provide software patches, i.e., they failed to work in realistic settings and generalize beyond the community benchmarks. For example, in the Functional Error track for Java, the best tool was able to correctly fix only out of previously unseen bugs (i.e., 10% fix rate) and many techniques were unable to provide any correct patches. The benchmarks had been curated with new and unforeseen repair tasks, and the tools had been evaluated in a uniform environment enforcing strict resource constraints.
@@
Machine Learning (ML) techniques, specifically Large Language Models (LLM), play a significant role in the SE and APR community’s most recent research efforts. Due to their pre-training on large amounts of open-source code and due to their enormous capacity, LLMs have demonstrated problem-solving skills over a wide range of tasks (Zhao et al., 2023). In APR, recent works (Xia and Zhang, 2022, 2024) have shown that LLMs can be used to generate repairs and build SE agents (Zhang et al., 2024b; Ruan et al., 2024; Yang et al., 2024), which attempt to mimic the work of software engineers. Nevertheless, most of the current LLM-based APR approaches use LLMs in a rather ad-hoc strategy (e.g., (Xia and Zhang, 2022; Xia et al., 2024b; Zhu et al., 2023; Hossain et al., 2024)), and while some existing works study the usage of LLMs in a more systematic way (e.g., (Huang et al., 2023; Silva et al., 2024; Jiang et al., 2023)), they all fall short on at least one of the following points: (i) the investigation is based on one specific LLM only, (ii) no/limited exploration of various prompting strategies/designs is conducted, in particular regarding repair ingredients (like information about failing test cases, bug reports/descriptions, error messages, etc.), (iii) a perfect Fault Localization (FL) is assumed to be given, which represents an unrealistic assumption in a practical setting, and (iv) only one specific benchmark with one programming language is covered, e.g., Defects4J (Just et al., 2014).
@@
To address the above issues and to provide a comprehensive overview of the current research landscape of APR, this work presents a systematic empirical evaluation with state-of-the-art closed and open LLMs on multiple benchmarks across programming languages. Hereby, we refer to open models as those with publicly available weights111Note that, according to our definition, open models may not necessarily come with open-source licenses, and that – unlike their weights – their training code and/or data may remain undisclosed., whereas closed LLMs are proprietary models whose weights, training data, and model architectures have not been publicly released. Particularly, we compare closed models like Claude 3.7 Sonnet, Gemini 1.5 Pro, OpenAI o3-mini, and OpenAI GPT-4o with open ones like Llama 3.3, CodeLlama, and DeepSeek Coder on benchmarks for Java (Defects4J (Just et al., 2014)), JavaScript (BugsJS (Gyimesi et al., 2021)), Python (BugsInPy (Widyasari et al., 2020)), and PHP (BugsPHP (Pramod et al., 2024)). In total, we generated and evaluated patches in four programming languages and fixed out of bugs.
@@
Based on our experiments, current LLMs do not generalize well across programming languages for the task of program repair. In particular, we observed that different LLMs perform well on different benchmarks. Similar to others (Parasaram et al., 2024), we observed that including information about the failing test case significantly boosted the repair performance for all models. We found LLMs to show good robustness when repairing bugs of higher complexity. However, notably, our comparison between perfect and automated FL techniques showed a drastic decline in repair capabilities if the buggy function cannot reliably be identified. Overall, our findings provide fundamental and comprehensive understanding of the current capabilities of LLMs with respect to the various aspects of APR, which is beneficial for our community (e.g., for designing future research efforts) and for practitioners (e.g., for selecting appropriate LLMs for their repair use case). In summary, we make the following core contributions:
@@
To the future APR researcher, our study shows that (1) fault localization is a potential bottleneck for APR pipelines, which suggests a strong focus on this aspect. (2) APR accuracy generalizes poorly, with different models reacting differently when varying programming language or prompt information; and (3) combining different LLMs improves results across languages and offers interesting directions for future work. We conclude that APR pipelines need to be benchmarked carefully and broadly. To support such a holistic evaluation of future progress in APR, we open-source our experimental setup and results for other researchers.
2. Background and Related Work
2.1. LLM based Program Repair
Numerous APR techniques have been studied in the last decade, ranging from the traditional search-based repair (Le Goues et al., 2012), constraint-based repair (Nguyen et al., 2013), template-based repair (Liu et al., 2019b), and learning-based repair (Long and Rinard, 2016). More recently, with the advancement of LLMs for code generation, various techniques (Xia and Zhang, 2022, 2024; Ye and Monperrus, 2024) have been explored by the research community to utilize LLMs for the program repair problem. LLM-based repair techniques transform the program repair problem into a code generation problem. These techniques can be broadly classified as model training (Silva et al., 2024; Xia et al., 2024b), zero-shot learning (Xia and Zhang, 2022) and few-shot learning (Tang et al., 2024; Jin et al., 2023; Yin et al., 2024). Silva et al. (Silva et al., 2024) explored how different input and output representations yield optimized results using LLMs for program repair tasks. Xia et al. (Xia and Zhang, 2024) studied the effect of test diagnostic feedback for LLM-driven program repair. Lin et al. (Lin et al., 2024) studied the impact of repair granularities encountered in program repair tasks. More specifically, they examined how expression-level and token-level repairs differ for different LLMs.
@@ 2.1. LLM based Program Repair
2.2. Empirical Studies of Program Repair
@@
Several empirical studies have been conducted to evaluate program repair techniques in their efficacy (Liu et al., 2020), performance (Zhong et al., 2023), trust (Noller et al., 2022), and overfitting (Durieux et al., 2019; Ye et al., 2019) aspects. Ye et al. (Ye et al., 2019) conducted an empirical study using 10 pre-LLM program repair tools for Java programs and found overfitting to the respective test suites to be an issue. Similarly, Durieux et al. (Durieux et al., 2019) conducted a large-scale evaluation using 11 APR tools to fix 2,141 Java programs in five different repair benchmarks. Their experiments showed that repair tools do not generalize across benchmarks, highlighting the problem of benchmark overfitting in repair techniques. Liu et al. (Liu et al., 2020) conducted another study on the efficiency of repair tools using 16 Java repair tools. Their study highlighted the performance inefficiencies in repair tools, which predominantly generated a large number of incorrect patches despite their overall repair capabilities. Noller et al. (Noller et al., 2022) performed a small-scale experiment using five program repair tools and 60 C programs. Their findings reported that different experiment configurations lead to different results, highlighting the importance of fair comparison under a uniform setup. A similar study was conducted by Zhong et al. (Zhong et al., 2023) for learning-based repair tools, highlighting the importance of a unified experimental setup with respect to the training data and evaluation metrics used for comparisons.
@@
More recently, Xia et al. (Xia et al., 2023) conducted an extensive evaluation using 9 LLMs and five repair benchmarks covering multiple languages under different repair configurations. Their findings highlighted the benefits of using LLMs for APR, with respect to the quality of the generated patches and the performance improvements compared to non-LLM-based APR tools. Several other studies (Jiang et al., 2023; Huang et al., 2023) investigated fine-tuned coding models for program repair using multiple benchmarks, showing they can outperform non-LLM-based repair tools. Ouyang et al. (Ouyang et al., 2024) studied 11 repair tools including traditional techniques, NMT-based, and LLM-based repair tools. Their findings further emphasized the benefits of using LLM-based repair tools, which showed less susceptibility to overfitting and better generalization across benchmarks. Unlike existing work, this study performs an extensive analysis using the most recent 13 LLMs including the most advanced reasoning models across various experiment configurations. We focus on the generalizability of LLMs across different languages, repair configurations and patch complexities. Additionally, we revisit the impact of accurate fault localization on program repair with respect to the usage of the most advanced reasoning LLMs.
3. Study Methodology
@@
The primary objective of this work is to explore the feasibility of APR using current LLMs. We first discuss the four research questions addressed in our study (Section 3.1), followed by the LLMs we evaluate (Section 3.2) and the benchmark datasets in different programming languages (Section 3.3). We outline further technical details – the evaluation protocol, technical setup and prompt design – in Sections 3.4 to 3.5.
3.1. Research Questions
@@
To answer this question, we apply LLM-based APR on repair benchmarks for Java, Python, PHP, and JavaScript and assess the quality of the generated patches. As part of this question, we also investigate whether one model appears superior across all programming languages, or whether different models work best on specific languages (which would suggest a language-specific model selection). To investigate this question, we use two prompt designs (see Section 3.5); (1) a basic prompt including only the code passage to be repaired, and (2) an extended prompt which also includes additional information from failing tests’ error messages.
@@
Traditionally, APR pipelines feature a fault localization (FL) step (in which a faulty code subsegment is identified) and a repair step (in which the subsegment is fixed). LLMs could either follow this approach and take localization hints from external systems such as FLACOCO (Silva et al., 2021) using spectrum-based fault localization (de Souza et al., 2016), or operate in an end-to-end fashion, fixing a given chunk of code without explicitly localizing faulty subsegments. So far, there are no conclusive research results as to what extent LLMs benefit from localization, with a lot of research assuming perfect localization to be given (Xia and Zhang, 2022, 2024; Hossain et al., 2024; Huang et al., 2023; Silva et al., 2024; Jiang et al., 2023), others focusing on end-to-end operation (Silva and Monperrus, 2024), and only few recent comparisons of both with inconclusive results so far (Silva et al., 2024; Xu et al., 2024). To this end, we investigate whether current LLMs benefit from localization signals at code line granularity.
@@
In our last question, we focus on the differences in APR accuracy between open LLMs and closed LLMs. While closed models have been found to offer superior performance on the majority of natural language understanding (NLU) tasks in the past (e.g., (Zhao et al., 2023)), open LLMs offer the benefit of being operational locally, thus allowing for better reproducibility of results and supporting technological sovereignty and democratization. Therefore, we track the progress of both categories of LLMs over time, to investigate whether open models are catching up with their closed counterparts.
3.2. Models
@@ 3.3. Benchmarks
Finally, the benchmarks should offer a sufficient number of bugs that are still reproducible today. As Zhu et al. (Zhu and Rubio-González, 2023) have shown, reproducibility is often a challenge in executable software defects datasets, as legacy versions of software artifacts are prone to breakage due to missing/deprecated dependencies or external resources.
Similar to prior studies (Xia and Zhang, 2022; Xia et al., 2023; Ouyang et al., 2024), we identified bugs of comparable complexity within the benchmarks to enable fair comparisons across programming languages. To do so, we evaluated single-function fixes within the benchmarks, meaning we considered only those bugs that could be resolved by modifications within a single function. Unlike many previous studies, however, we did not restrict our analysis to single-line (Liu et al., 2019b; Kolak et al., 2022) or single-hunk fixes (Xia and Zhang, 2022), that is, bugs that can be resolved by modifying a single contiguous piece of code.
@@
Defects4J v3.0 (Just et al., 2014)
This curated collection of reproducible bugs, supported by a robust infrastructure, has been widely adopted in software engineering research. Most prior studies on LLM-based APR were performed on Defects4J (Just et al., 2014), thus enabling a direct comparison with prior work. It contains 854 bugs from 17 open-source Java projects. For our first research question, we used the 264 single-hunk bugs that could be reproduced.
BugsInPy (Widyasari et al., 2020)
This benchmark dataset comprises 493 real bugs from 17 open-source Python projects across diverse domains, including machine learning, developer tools, scientific computing, and web frameworks. Unlike Defects4J, which distinguishes between active and deprecated bugs (i.e. those no longer reproducible with current Java versions), BugsInPy does not track the reproducibility status of its bugs. Aguilar et al. (Aguilar et al., 2023) evaluated the reproducibility of the dataset three years after its initial release and found that only of the expected results could be reproduced in the unmodified BugsInPy dataset. Of these 337 reproducible bugs, 82 were single-hunk bugs, which we used in our first experiment.
BugsJS (Gyimesi et al., 2021)
@@
BugsPHP (Pramod et al., 2024)
BugsPHP is a benchmark dataset of PHP bugs in real-world applications consisting of a training and test dataset. We use BugsPHP’s test split, which contains 513 bug-fixing commits from 15 popular open-source PHP projects covering the time period from Jan 2020 to Mar 2023. Each commit is equipped with at least one relevant failing test case. For our initial experiment, we identified 129 reproducible single-hunk bugs.
| Benchmark | all bugs | single-file | single-function | single-hunk | single-line |